# Contest for system observability as an infinitely repeated game

### v2

@@ 4.2 Game based on two-step optimality
$BR2X^{a}=\arg\max_{U_{a}\in\mathbb{R}^{k\times n}}[\text{MGM}(N+B_{a}^{\prime}U_{a})].$ (31)
$BR2X^{a}$ is not the best response set for the attacker considering two-step optimality, since it does not guarantee that the value function of the current epoch is maximized.
Because $\max_{U_{a}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=\operatorname{dim}\mathcal{V}^{*}(H)$, the best response set for the defender following two-step optimality is
$\begin{split}BR2^{d}=&arg\min_{H\in BR1^{d}}\max_{U_{a}\in\mathbb{R}^{k\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)\\ =&arg\min_{H\in BR1^{d}}\operatorname{dim\,}\mathcal{V}^{*}(H).\end{split}$ (32)
Define set $BR2X^{d}$ as
$BR2X^{d}=\arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H).$ (33)
$BR2X^{d}$ does not ensure the value function of the current epoch is minimized; hence, it is not the best response set for the defender considering two-step optimality.
Figure 4: Sequence of actions following two-step optimality.
Remark 4.13.
Actually, $BR2X^{d}$ and $BR2X^{a}$ are Stackelberg game equilibrium in two different scenarios. In a Stackelberg game, the leader possesses knowledge of the follower’s utility function and takes the initial action, taking into account how the follower is likely to respond. The follower, on the other hand, has the ability to observe the leader’s action and change its own action accordingly. In scenario 1, the defender is the leader and the attacker is the follower. The best response of the follower is $BR1^{a}$ and the Stackelberg equilibrium action of the leader is $BR2X^{d}$. In scenario 2, the attacker is the leader and the defender is the follower. The best response of the follower is $BR1^{d}$ and the Stackelberg equilibrium action of the leader is $BR2X^{a}$.
Fig. 4 shows the sequence of actions. Then we define the Nash equilibrium when two players consider two-step optimality.
Definition 3.
The strategy profile $(U_{a}^{*},H^{*})$ is said to be a Nash equilibrium of the two-step optimal game if $U_{a}^{*}\in BR2^{a},H^{*}\in BR2^{d}$.
In order to get the condition for Nash equilibrium, we give the following lemma.
Lemma 3.
For the same $N\in\mathbb{R}^{n\times n},B_{a}^{\prime}\in\mathbb{R}^{n\times k}$, $\forall U_{a}\in\mathbb{R}^{k\times n}$, there is $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)\geq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H)$.
Proof 4.14.
$\forall\,H_{0}\in\mathbb{R}^{m\times n}$, it is apparent that the maximal controlled invariant subspace in Ker $H_{0}$ is larger than $(N+B_{a}^{\prime}U_{a})$-invariant subspace in Ker $H_{0}$, i.e., $\operatorname{dim\,}\mathcal{V}^{*}(H_{0})\geq\operatorname{dim\,Ker}\Omega(U_{a},H_{0}).$ Thus $\forall H_{1}\in arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)$, $\operatorname{dim\,}\mathcal{V}^{*}(H_{1})\geq\operatorname{dim\,Ker}\Omega(U_{a},H_{1})$. And $\forall H_{2}\in\arg\min_{{H}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}$ $\Omega(U_{a},H)$, there is $\operatorname{dim\,Ker}\Omega(U_{a},H_{1})\geq\operatorname{dim\,Ker}\Omega(U_{a},H_{2})$. Thus $\operatorname{dim\,}\mathcal{V}^{*}(H_{1})\!\geq\!\operatorname{dim\,Ker}\Omega(U_{a},H_{2})$.
Based on this lemma, it is found that the Nash equilibrium condition in Theorem 4.5 is also valid when two players consider two-step optimality. Although the equilibrium conditions for the player considering one-step or two-step optimality are the same, the mathematical reasoning is different.
Theorem 4.15.
The strategy profile $(U_{a}^{*},H^{*})$ is a Nash equilibrium when two players consider two-step optimality if and only if, $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H)=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})=\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})$.
Proof 4.16.
(Sufficiency) According to Theorem 4.5, $H^{*}\in BR1^{d}$ and $U_{a}^{*}\in BR1^{a}$. We need to further prove $H^{*}\in BR2^{d}$ and $U_{a}^{*}\in BR2^{a}$. $\forall H\in BR1^{d}$, $\operatorname{dim\,}\mathcal{V}^{*}(H)\geq\operatorname{dim\,Ker}\Omega(U_{a}^{*},H)=\gamma,$ where $\gamma$ is a constant for different $H\in BR1^{d}$. Because $\exists H^{*}\in BR1^{d}$, $\operatorname{dim\,}\mathcal{V}^{*}(H^{*})$ $=\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H^{*})=\gamma$, $\operatorname{dim\,}\mathcal{V}^{*}(H^{*})$ reaches the lower bound. Thus $H^{*}\in BR2^{d}$. Define $F_{1}(U_{a})=\min_{H\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega(U_{a},H),$ then the upper bound of $F_{1}(U_{a})$ is $\overline{F_{1}}\!=\!\max_{U_{a}\in\mathbb{R}^{k\times n}}\!\min_{H\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}$ $\Omega(U_{a},H).$ According to Lemma 3, $\overline{F_{1}}\leq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}$ $\mathcal{V}^{*}(H)$ $\leq\operatorname{dim\,}\mathcal{V}^{*}(H\in\mathbb{R}^{m\times n})$. Since $F_{1}(U_{a}^{*})=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})$, there must be $F_{1}(U_{a}^{*})=\overline{F_{1}}=\min_{H\in\mathbb{R}^{m\times n}}$ $\operatorname{dim\,}\mathcal{V}^{*}(H)=\operatorname{dim\,}\mathcal{V}^{*}(H^{*}).$ Because $U_{a}^{*}$ lets $F_{1}$ reaches the upper bound, $U_{a}^{*}\in BR2^{a}$. Thus $(U_{a}^{*},H^{*})$ is a Nash equilibrium under the game framework of considering two-step optimality.
(Necessity) Since $BR2^{d}\in BR1^{d}$ and $BR2^{a}\in BR1^{a}$, the rest proof is the same as Theorem 4.5.
Although the best response sets for two-step optimality are much smaller than that for one-step optimality, they still are not a singleton. Therefore, we extend Assumption 1 and 2 to the game framework considering two-step optimality. Then we can easily find that corollaries 4.8-4.11 also hold in game following two-step optimality.
4.3 General equilibrium analysis
This subsection gives conclusions that are consistent with both one-step and two-step optimality game. Firstly, we introduce a special equilibrium condition.
Theorem 4.17.
For $N\in\mathbb{R}^{n\times n},B_{a}^{\prime}\in\mathbb{R}^{n\times k}$, if $\exists$ $H^{*}\in\mathbb{R}^{m\times n}$ satisfies $(1)$ or $(2)$:
[…]
then $\operatorname{dim}$ $\mathcal{V}^{*}(H^{*})=0.$ And $\forall U_{a}\in\mathbb{R}^{k\times n}$ the strategy profile $(U_{a},H^{*})$ is an equilibrium.
Proof 4.18.
$(1)\text{ When }(n-m)\geq k.\text{ Since }\operatorname{Im}B_{a}^{\prime}\subseteq\operatorname{Ker}H^{*},H^{*}B_{a}^{\prime}=0.\text{ Then }\forall v\neq 0\in\operatorname{Ker}H^{*},\forall U_{a},H^{*}(N+B_{a}^{\prime}U_{a})v=H^{*}Nv\neq 0.$ Thus $(N+B_{a}^{\prime}U_{a})v\notin\operatorname{Ker}H^{*}$, which means controlled invariant subspace contained in $\operatorname{Ker}H^{*}$ is 0. Thus dim $\mathcal{V}^{*}(H^{*})=0.$
[…]
When dim $\mathcal{V}^{*}(H^{*})=0$, $\forall U_{a}\in\mathbb{R}^{k\times n}$, $\operatorname{dim\,Ker}\Omega$ $(U_{a},H^{*})=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})=0$. And it is apparent that $H^{*}\in\arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega$ $\,(U_{a},H)$. Thus $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=\operatorname{dim\,}\mathcal{V}^{*}$ $(H^{*})=\operatorname{dim\,Ker}\Omega(U_{a},H^{*})$. According to Theorems 1 and 2, $(U_{a},H^{*}),\forall U_{a}\in\mathbb{R}^{k\times n}$ is an equilibrium whenever two players consider one-step or two-step optimality.
In this scenario, the defender can identify such an $H^{*}$ that makes $\mathcal{V}^{*}$ equal to zero. Then the attacker is unable to alter observability of the system any more and a trivial result is obtained. To avoid this special equilibrium case, set the game problem based on $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}>0.$ Then we want to give a condition for this setting.
Theorem 4.19.
For system (9), $B_{a}^{\prime}\in\mathbb{R}^{n\times k}$ and $H\in\mathbb{R}^{m\times n}$ are of full rank, if
$\max\{k,\text{MGM}({N}_{\bar{r}}))\}>m,$ (34)
where $\text{MGM}({N}_{\bar{r}})$ is the maximum geometric multiplicity associated with the eigenvalues of ${N}_{\bar{r}},$ ${N}_{\bar{r}}$ is the uncontrollable part for system matrix $N$, then $\min_{H\in\mathbb{R}^{m\times n}}$ $\operatorname{dim\,}\mathcal{V}^{*}>0.$
@@ 4.3 General equilibrium analysis
Let the subscript $r$ represent the controllable part and the subscript $\bar{r}$ represent the uncontrollable part. The maximum controlled invariant subspace $\mathcal{V}^{*}$ can be divided into controllable part $\mathcal{V}^{*}_{r}$ and uncontrollable part $\mathcal{V}^{*}_{\bar{r}}$, where $\mathcal{V}^{*}_{r}$ is determined by $(N_{r},B^{\prime}_{ar},H_{r})$ and $\mathcal{V}^{*}_{\bar{r}}$ is determined by $(N_{\bar{r}},0,H_{\bar{r}})$.
(1) If $k>\text{MGM}({N}_{\bar{r}})$, then $k>m$. For controllable part $(N_{r},B^{\prime}_{ar},H_{r})$, the geometric multiplicity of $(N_{r}+B^{\prime}_{ar}U_{ar})^{\prime}s$ eigenvalue $\lambda$ is $\{n-\text{rank}[(N_{r}+B^{\prime}_{ar}U_{ar})-\lambda I]\}.$ According to PBH controllability criterion, for the controllable part, $\forall\lambda\in\sigma(N_{r}+B^{\prime}_{ar}U_{ar})$, rank$[(N_{r}\!+\!B^{\prime}_{ar}U_{ar}\!-\!\lambda I)\,\,\,B^{\prime}_{ar}]=n$. Thus rank$[(N_{r}\!+\!B^{\prime}_{ar}U_{ar}\!-\!\lambda I)]\geq(n-k)$ and $\text{MGM}(N_{r}\!+\!B^{\prime}_{ar}U_{ar})=k.$ If $k>m$, i.e., $\text{MGM}(N_{r}\!+\!B^{\prime}_{ar}U_{ar})>m$, according to Corollary 3.3, =$\mathcal{I}=\{i\mid\alpha_{i}>m\}\neq\emptyset,\underset{{H_{r}}\in\mathbb{R}^{m\times n}}{\min}\operatorname{dim\,Ker}\Omega(U_{ar},H_{r})=n-\underset{{q}\in\mathcal{I}}{\min}(\sum_{p=1}^{m}r_{qj_{p}})>0.$ According to Lemma 3, $\min_{H_{r}\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}_{r}^{*}\geq\min_{H_{r}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{ar},H_{r})>0$. Thus $\min_{{H}_{r}}\operatorname{dim\,}$ $\mathcal{V}_{r}^{*}>0$.
(2) If $\text{MGM}({N}_{\overline{r}})>k$, then $\text{MGM}({N}_{\overline{r}})>m$. For uncontrollable subspace, $\mathcal{V}^{*}_{\bar{r}}=\operatorname{Ker}\Omega(N_{\bar{r}},0,H_{\bar{r}})$. If $\text{MGM}({N}_{\overline{r}})>m$, extending Lemma 2 to the dual system, we have $\min_{{H}_{\overline{\mathrm{r}}}}\operatorname{dim\,Ker}\Omega>0$. Thus $\min_{{H}_{\overline{\mathrm{r}}}}\operatorname{dim\,}\mathcal{V}_{\bar{r}}^{*}>0$.
In (1) or (2), $\operatorname{dim\,}\mathcal{V}^{*}=\operatorname{dim\,}\mathcal{V}_{r}^{*}+\operatorname{dim\,}\mathcal{V}_{\bar{r}}^{*}>0.$
Under the conditions of $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}>0$, we obtain the following theorem.
Theorem 4.21.
For $\mathcal{V}_{1}^{*}=\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}$, if $\operatorname{dim\,}\mathcal{V}_{1}^{*}>0$ and $U_{a}^{*}\notin\mathcal{F}(\mathcal{V}_{1}^{*})$, then $\forall H\in\mathbb{R}^{m\times n},(U_{a}^{*},H)$ is not a Nash equilibrium.
Proof 4.22.
We prove it by the contrapositive: if $\exists H^{*}$ which makes $(U_{a}^{*},H^{*})$ a Nash equilibrium, then $U_{a}^{*}\in\mathcal{F}(\mathcal{V}_{1}^{*})$, where $\mathcal{V}_{1}^{*}=\min_{H\in\mathbb{R}^{m\times n}}\mathcal{V}^{*}(H)$. According to Lemma 3, $\forall U_{a}\in\mathbb{R}^{k\times n}$, $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H)\leq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)$ $\leq\operatorname{dim\,}\mathcal{V}^{*}(H\in\mathbb{R}^{m\times n})$. According to Theorems 1 and 2, if $(U_{a}^{*},H^{*})$ is a Nash equilibrium, $(U_{a}^{*},H^{*})$ satisfies $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,$ $(U_{a}^{*},H)=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})$. Thus there is $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}$ $\Omega(U_{a}^{*},H)=\min_{H\in\mathbb{R}^{m\times n}}$ $\operatorname{dim\,}\mathcal{V}^{*}(H)=\operatorname{dim\,}\mathcal{V}^{*}(H^{*}).$ Therefore, we have $H^{*}\in arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}$ $\mathcal{V}^{*}(H)$ and $\mathcal{V}^{*}(H^{*})=\mathcal{V}_{1}^{*}.$ Because $\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})=\operatorname{dim\,}\mathcal{V}_{1}^{*}$ according to Theorems 1 and 2, $U_{a}^{*}\in\mathcal{F}(\mathcal{V}_{1}^{*})$. Thus the contrapositive has been proven.
4.4 Discussions
According to Theorems 1-3, we find that the defender dominates the game to achieve equilibrium. On the one hand, the defender can achieve equilibrium of the game by considering two-step optimality. This is because for fixed $U_{a}$, $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=\operatorname{dim\,Ker}\Omega(U_{a},H\in BR1^{d})\leq\operatorname{dim\,}\mathcal{V}^{*}(H\in BR2^{d})\leq\operatorname{dim\,}\mathcal{V}^{*}(H\in BR1^{d})$. $H\in BR2^{d}$ makes $\operatorname{dim\,}\mathcal{V}^{*}$ closer to $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)$ compared to $H\in BR1^{d}$. If they are equal, the condition for Nash equilibrium is satisfied. As for the attacker, it plays little role in generating equilibrium. Because for fixed $H_{1}$, $U_{a}^{*}\in BR1^{a}$ ((((or $U_{a}^{*}\in BR2^{a})$, there must be $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H)\leq\operatorname{dim\,}\mathcal{V}^{*}(H_{1})=\operatorname{dim\,Ker}\Omega(U_{a}^{*},H_{1})$. Because the relation between $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H)$ and $\operatorname{dim\,}\mathcal{V}^{*}(H_{1})$ is unrelated to $U_{a}^{*}$, the attacker cannot make the game converge to equilibrium. On the other hand, Theorem 4.17 gives a case when the action of the defender and system matrices satisfy a specific condition, the game invariably remains in equilibrium regardless of the attacker’s action.
Considering Theorem 4.21, we find that the attacker can break the equilibrium regardless of the defender’s action if $U_{a}^{*}\notin\mathcal{F}(\mathcal{V}_{1}^{*}),$ where $\mathcal{V}_{1}^{*}$ is the smallest controlled invariant subspace. But in this case, the attacker does not maximize the current value function, forsaking the current gain. In fact, Nash equilibrium is sometimes disadvantageous to the attacker because the value function remains at a low value that the attacker cannot change. In this case, the attacker can break the equilibrium by choosing $U_{a}^{*}\notin\mathcal{F}(\mathcal{V}^{*})$ and entice the defender to choose a lower value function action in the next epoch. However, the corresponding $\mathcal{V}^{*}$ value might be much higher, leading to the attacker achieving a much higher value function in the subsequent second epoch and in the future epochs. An example is given in Section 5 to illustrate this case.
To conclude, because the best response sets considering one-step optimality or two-step optimality are not single-valued maps, different choices in the best response sets will lead to different game results. Nevertheless, three rules of the game results can be summarized:
(1) The conditions for Nash equilibrium are the same whether players consider one-step optimality or two-step optimality.
(2) If both players are greedy, i.e., their actions ensure the optimum for the current epoch, the defender can find the Nash equilibrium following two-step optimality and has a dominant role in determining whether the game can reach an equilibrium.
(3) The attacker, by sacrificing current gains, can break the Nash equilibrium, enticing the defender to choose a different action associated with a lower value function and making the game become another pattern. Then the attacker may achieve a much higher value function in the future epochs.
In addition, this work only considers one-step optimality and two-step optimality, but does not consider three-step optimality or more steps, because it is the same player changing actions in the first step and in the third step. The game scenario in the first epoch is the same as that in the third epoch. So there is no need to consider higher steps.
@@ 5 Illustrative examples
$\hat{y}$ $=Hz,$
with $N={\left[\begin{array}[]{lllll}0.3&0&0&0&0\\ 0&0.3&0&0&0\\ 0&0&0.3&0&0\\ 0&0&0&0.1&0\\ 0&0&0&0&0.2\end{array}\right]},B_{a}^{\prime}=\left[\begin{array}[]{lllll}0&0&1&0&1\end{array}\right]^{\top}$. The attacker controls $U_{a}\in\mathbb{R}^{1\times 5}$ to maximize the unobservable subspace while the defender controls $H\in\mathbb{R}^{2\times 5}$ to minimize the unobservable subspace, forming an infinitely repeated game. In epoch 1, $U_{a}=\left[\begin{array}[]{lllll}0&0&0&0&0\end{array}\right]$. The examples in this section are periodic. Although only the outcomes of 60 epochs are shown in the following figures, they are sufficient to represent the scenarios over an infinite time span.
Table 2: Game settings for Cases 1-6
@@
Case 1 Algorithm 1. Algorithm 2. One-step optimality
Case 2 $U_{a}\!=\!\left[\begin{array}[]{lllll}0&0&0&0.2&0.1\end{array}\right]$ instead of $\left[\begin{array}[]{lllll}\!-0.1&0&0&0.1&0\end{array}\right]$ when $H\!=\!\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right],$ and other actions based on Algorithm 1. Algorithm 2.
Case 3 Algorithm 1. $H\!=\!\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&0&1&0&0\end{array}\right]$ instead of $\!\!\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right]$ when $U_{a}\!=\!\left[\begin{array}[]{lllll}0&0&0&0&0\end{array}\right],$ and other actions follow Algorithm 2.
Case 4 Before epoch 20: Algorithm 1; after 20 epochs: two-step optimality. Before epoch 40: Algorithm 2; after 40 epochs: two-step optimality. One-step & two-step optimality
Case 5 Before epoch 40: Algorithm 1; after 40 epochs: two-step optimality. Before epoch 20: Algorithm 2; after 20 epochs: two-step optimality.
Case 6 One-step optimality, only with $U_{a}\notin\mathcal{F}(\mathcal{V}^{*})$ in epoch 40. One-step optimality. One-step optimality & attacker not greedy
Table 3: Actions and average value functions of players for Cases 1-3.
@@
$\bar{\Phi}$ 1.5 2 1
Cases 1-3: let players have different choices in the one-step optimal best response sets. In Case 1, the attacker chooses $U_{a}\!=\!\left[\begin{array}[]{lllll}\!-0.1&0&0&0.1&0\end{array}\right]$ when $H\!=\!\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right]$ according to Algorithm 1, which uses pseudo inverse to get $U_{a}$ with the minimum modulus length; the defender chooses $H=\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right]$ when $U_{a}\!=\!\left[\begin{array}[]{lllll}0&0&0&0&0\end{array}\right]$ according to Algorithm 2. While in Cases 2-3, players have different choices compared to Case 1.
Figure 5: Evolution of value function for Cases 1-3. (a) Case 1: the attacker uses Algorithm 1 and the defender uses Algorithm 2; (b) Case 2: the attacker changes strategy preference and other actions match Case 1; (c) Case 3: the defender changes strategy preference and other actions match Case 1.
Fig. 5 shows the evolution of unobservable subspace dimension for Cases 1-3, which are totally different in different cases. More specifically, the actions of players $U_{a}$ and $H$ evolve in different loops, which are summarized in Table 3. The average value functions $\bar{\Phi}$ are $1.5,2,1$ respectively for Cases 1-3.
The reasons for different utilities in different cases are as follows: in Case 1, $U_{a}\!=\!\left[\begin{array}[]{lllll}\!-0.1&0&0&0.1&0\end{array}\right]\in BR1^{a}$, but $U_{a}\notin BR2^{a};H=\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right]\in BR1^{d}$, but $H\notin BR2^{d}$. In Case 2, $U_{a}=\left[\begin{array}[]{lllll}0&0&0&0.2&0.1\end{array}\right]\in BR2^{a}.$ In Case 3, $H=\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&0&1&0&0\end{array}\right]\in BR2^{d}$. As expected, when the action of the player falls exactly within its two-step optimal set, it can make the average value function more favorable for itself.
Cases 1-2 illustrate the results of Corollary 4.10 and 4.11. In Case 3, the value function and actions $U_{a}$ and $H$ are locked. Such $(U_{a},H)$ is a Nash equilibrium, which satisfies $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=\operatorname{dim\,}\mathcal{V}^{*}(H)$. Case 3 illustrates the results of Theorem 4.5 and Corollary 4.8.
Figure 6: Game results for Cases 4-5. (a) Evolution of value function for Case 4; (b) Evolution of $\mathcal{V}^{*}$ space dimension for Case 4; (c) Evolution of value function for Case 5; (d) Evolution of $\mathcal{V}^{*}$ space dimension for Case 5.
Figure 7: Results of Case 6: game reaches equilibrium before epoch 40 and after epoch 40, the defender choose $U_{a}\notin BR1^{a}$. (a) Evolution of value function; (b) Evolution of $\mathcal{V}^{*}$ space dimension.

### v3

@@ 4.4 Equilibrium characterization
$U_{a}^{*}\notin\mathcal{F}(\mathcal{V}^{*}(H^{\prime})),\text{where }H^{\prime}\in\arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H).$ (35)
Then, for every $H\in\mathbb{R}^{m\times n}$, the profile $(U_{a}^{*},H)$ is not a Nash equilibrium.
@@
When $\min_{H}\operatorname{dim}\mathcal{V}^{*}(H)=0$, there is $\mathcal{V}^{*}(H^{\prime})=\{0\}$. By the definition of $\mathcal{F}$ we have ($\mathcal{F}(\{0\})=\mathbb{R}^{k\times n}$, so no attacker strategy can satisfy (35). When $\min_{H}\operatorname{dim}\mathcal{V}^{*}(H)>0$, $\mathcal{V}^{*}(H^{\prime})$ is a non-trivial proper subspace of $\mathbb{R}^{m}$, implying $\mathcal{F}(\mathcal{V}^{*}(H^{\prime}))\subsetneq\mathbb{R}$. Hence $\mathbb{R}\setminus\mathcal{F}(\mathcal{V}^{*}(H^{\prime}))\neq\varnothing$; choose any $U_{a}^{*}$ in this complement. Such a $U_{a}^{*}$ satisfies (35) and therefore violates the necessary condition (33). By Theorem 4.16, $(U_{a}^{*},H)$ fails to satisfy the necessary condition for a Nash equilibrium for all $H\in\mathbb{R}^{m\times n}$. Hence $(U_{a}^{*},H)$ cannot be a Nash equilibrium, which completes the proof.
By Corollary 4.19, the attacker can break any candidate equilibrium by selecting $U_{a}^{*}\notin\mathcal{F}\bigl{(}\mathcal{V}^{*}(H^{\prime})\bigr{)},$ where $\mathcal{V}^{*}(H^{\prime})$ is the minimal controlled invariant subspace. Although this choice forgoes the one–period optimum, it forces the defender to deviate in the next epoch. That deviation enlarges the invariant subspace and raises the attainable value function. The attacker can then switch to an action compatible with the new subspace, earning a strictly higher utility from the second epoch on and maintaining that advantage thereafter. Section 5 presents an example that illustrates this case.
4.5 Summary
To conclude, this chapter analyze the one-step optimality and two-step optimality perspectives and three key insights of the observability-adversarial game can be summarized:
(1) In Subsections 4.1–4.2, Theorems 4.2 and 4.7 show that both the one-step and two-step formulations share a unified necessary-and-sufficient test for a Nash equilibrium, turning different planning horizons into a single easy check. Although the two-step best response sets lie strictly inside the one-step best response sets, the equilibrium reached under either horizon is identical.
(2) In subsection 4.3, as long as the defender selects an $H^{*}$ that satisfies the defender-dominated Nash equilibrium in Theorem 4.10, the controlled invariant subspace can be collapsed to zero, i.e. $\mathcal{V}^{*}(H^{*})=\{0\}$. According to Theorem 4.14, if in some round the profile $(U_{a},H)$ is a Nash equilibrium, then all subsequent rounds are locked at the same strategy pair (lock mode) and the value function remains fixed, leaving the attacker no further leverage to decrease observability. Hence, the defender not only terminates strategy evolution but also keeps system observability permanently at the most favorable level for the defender.
(3) In subsection 4.4, Theorem 4.16 states the necessary condition for Nash equilibrium. The attacker can purposely choose a $U_{a}^{*}$ violating this condition, sacrificing the immediate best response and thus invalidating the existing locked equilibrium. Once the equilibrium is broken, the strategy trajectory follows Theorem 4.14 into the Loop mode: both strategies and the value function oscillate on a finite cycle. During such a cycle, the controlled invariant subspace can enlarge and system observability may further degrade, so that by accepting a delayed utility the attacker potentially secures a higher long-term value and creates opportunities for deeper penetration.
@@ 5 Illustrative examples
$\widehat{y}$ $=Hz,$
with $N={\left[\begin{array}[]{lllll}0.3&0&0&0&0\\ 0&0.3&0&0&0\\ 0&0&0.3&0&0\\ 0&0&0&0.1&0\\ 0&0&0&0&0.2\end{array}\right]},B_{a}^{\prime}=\left[\begin{array}[]{lllll}0&0&1&0&1\end{array}\right]^{\top}$. Initialize with $U_{a}=\begin{bmatrix}0&0&0&0&0\end{bmatrix}$. The defender then selects $H\in\mathbb{R}^{2\times 5}$ to minimize the dimension of the unobservable subspace. Next, the attacker updates $U_{a}\in\mathbb{R}^{1\times 5}$ to maximize that dimension. This sequence of moves repeats and yields a repeated game. Although only the outcomes of 60 epochs are shown in the following figures, the cyclic nature of the results makes these snapshots sufficient to characterize the behavior over an infinite time horizon.
Table 2     Game settings for Cases 1–6
@@
Case 1 Algorithm 1. Algorithm 2. One‐step optimality
Case 2 $U_{a}\!=\!\bigl{[}\begin{array}[]{lllll}0&0&0&0.2&0.1\end{array}\bigr{]}$ instead of $\bigl{[}\begin{array}[]{lllll}-0.1&0&0&0.1&0\end{array}\bigr{]}$ when $H\!=\!\bigl{[}\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\bigr{]},$ and other actions based on Algorithm 1. Algorithm 2.
Case 3 Algorithm 1. $H\!=\!\bigl{[}\begin{array}[]{lllll}1&0&0&1&1\\ 0&0&1&0&0\end{array}\bigr{]}$ instead of $\bigl{[}\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\bigr{]}$ when $U_{a}\!=\!\bigl{[}\begin{array}[]{lllll}0&0&0&0&0\end{array}\bigr{]},$ and other actions follow Algorithm 2.
Case 4 Before epoch 20: Algorithm 1; after 20 epochs: two‐step optimality. Before epoch 40: Algorithm 2; after 40 epochs: two‐step optimality. One‐step & two‐step optimality
Case 5 Before epoch 40: Algorithm 1; after 40 epochs: two‐step optimality. Before epoch 20: Algorithm 2; after 20 epochs: two‐step optimality.
Case 6 One‐step optimality, only with $U_{a}\notin\mathcal{F}(\mathcal{V}^{*})$ in epoch 40. One‐step optimality. One‐step optimality & attacker not greedy
Cases 1-3: let players have different choices in the one-step optimal best response sets. In Case 1, the attacker chooses $U_{a}\!=\!\left[\begin{array}[]{lllll}\!-0.1&0&0&0.1&0\end{array}\right]$ when $H\!=\!\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right]$ according to Algorithm 1, which uses pseudo inverse to get $U_{a}$ with the minimum modulus length; the defender chooses $H=\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right]$ when $U_{a}\!=\!\left[\begin{array}[]{lllll}0&0&0&0&0\end{array}\right]$ according to Algorithm 2. While in Cases 2-3, players have different choices compared to Case 1. Figure 5 shows the evolution of unobservable subspace dimension for Cases 1-3, which are different for different cases. The specific actions of players $U_{a}$ and $H$ for cases 1-3 are summarized in Table 3. In cases 1-2, the actions $U_{a}$ and $H$ evolve in different loops, which illustrate the results of Theorem 4.14. In Case 3, the actions $U_{a}^{*}$ and $H^{*}$ keep unchanged, which is the candidate equilibrium. We then verify that $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a}^{*},H)\;=\;\dim\mathcal{V}^{*}(H^{*})$, thereby confirming the validity of Theorem 4.2 and Theorem 4.14.
Figure 5  Evolution of $\operatorname{dim\,Ker}\Omega$ for Cases 1-3. (a) Case 1: the attacker uses Algorithm 1 and the defender uses Algorithm 2; (b) Case 2: the attacker changes strategy preference and other actions match Case 1; (c) Case 3: the defender changes strategy preference and other actions match Case 1.
Table 3   Actions evolution of players for Cases 1-3.